### v1

# Boundary-Aware Adversarial Filtering for Reliable Diagnosis under Extreme Class Imbalance

Abstract
We study classification under extreme class imbalance where recall and calibration are both critical: for example, in medical diagnosis scenarios. We propose AF–SMOTE, a mathematically motivated augmentation framework that first synthesizes minority points and then filters them by an adversarial discriminator and a boundary utility model. We prove that, under mild assumptions on the decision boundary smoothness and class-conditional densities, our filtering step monotonically improves a surrogate of Fβ (for $\beta\!\geq\!1$) while not inflating Brier score. On MIMIC-IV proxy label prediction and canonical fraud detection benchmarks, AF–SMOTE attains higher recall and average precision than strong oversampling baselines (SMOTE, ADASYN, Borderline-SMOTE, SVM-SMOTE), and yields the best calibration. We further validate these gains across multiple additional datasets beyond MIMIC-IV. Our successful application of AF–SMOTE to a healthcare dataset using a proxy label demonstrates in a disease-agnostic way its practical value in clinical situations, where missing true positive cases in rare diseases can have severe consequences.
Index Terms—  Imbalanced learning, SMOTE, synthetic oversampling, adversarial filtering, medical diagnosis, calibration
1 Introduction
Fig. 1: AF–SMOTE Architecture. The Backbone synthesizes minority candidates via SMOTE and evaluates them with multi-branch heads: Realism (discriminator), Boundary utility (margin/probability), Uncertainty, and Density/outlier metrics. Scores are fused with learned weights and diversity regularization; Top-$K$ selection forms an augmented training set. The ADAPT Controller handles high-dimensional cases via PCA projection and gates hyperparameters, while the Focus Loop provides feedback to concentrate synthetic samples near under-represented boundary regions.
Extreme class imbalance ($\pi_{1}\!\ll\!1$) yields high apparent accuracy yet poor detection of positives; in clinical prediction this makes recall critical. Classical oversampling (e.g., SMOTE [chawla2002smote]) can harm calibration and fails to target boundary-relevant regions. We ask: where should synthetic points be placed and which should be kept? AF–SMOTE answers by generating candidates then filtering by realism (discriminator) and utility (boundary proximity), yielding a simple, analyzable rule with consistent empirical gains.
2 Notation & Assumptions
We denote by $\mathcal{D}$ the data distribution, with minority prior $\pi_{1}\!=\!P(y{=}1)\!\ll\!1$. Let $\hat{p}(x)=P_{\theta}(y{=}1\mid x)$ be a calibrated scorer and $t$ a decision threshold. The Bayes decision boundary is $\partial\mathcal{R}$, and $d(x,\partial\mathcal{R})$ is a differentiable surrogate of the distance to this boundary. We write $\rho$ for the reach of the minority manifold, $L$ the local Lipschitz constant of $\hat{p}$ near $\partial\mathcal{R}$, $\varepsilon$ the discriminator’s false-positive bound on real minority, $\lambda\in[0,1]$ the realism/utility weight, $\tau$ the selection threshold, and $p_{0}$ a precision floor. The combined score is $S(x)=\lambda s_{\text{util}}(x)+(1-\lambda)s_{\text{real}}(x)$.
Symbols.
$\hat{p}(x)$ calibrated probability scorer
$t$ decision threshold for positives
$\partial\mathcal{R}$ Bayes decision boundary
$d(x,\partial\mathcal{R})$ boundary-distance surrogate
$\rho$ reach of minority manifold
$L$ local Lipschitz constant of $\hat{p}$ near $\partial\mathcal{R}$
$\varepsilon$ discriminator false-positive bound on real minority
$\lambda,\tau$ realism/utility weight and selection threshold
$p_{0}$ target precision floor
$S(x)$ convex combination score   $\lambda s_{\text{util}}+(1-\lambda)s_{\text{real}}$
Assumptions.
  • (A1)
    Local Lipschitz. $\hat{p}$ is $L$-Lipschitz in a $\delta$-neighborhood of $\partial\mathcal{R}$.
  • (A2)
    Minority manifold & reach. Real minority support lies in a $C^{2}$ manifold $\mathcal{M}_{1}$ with reach $\rho$; SMOTE-like candidates remain in its tubular neighborhood.
  • (A3)
    Discriminator FP bound. For $x\sim P(\cdot\mid y{=}1)$, $\Pr[s_{\text{real}}(\tilde{x})\geq\eta\,\&\,\tilde{x}\notin\text{real minority}]\leq\varepsilon$.
  • (A4)
    Precision floor. Choose $\tau$ and then a threshold $t$ to satisfy $\mathrm{Prec}\geq p_{0}$ on validation.
  • (A5)
    Boundary-surrogate consistency. Higher $s_{\text{util}}$ implies no worse boundary distance up to a constant shift; local changes in $\hat{p}$ are bounded by distance to a boundary projection.
Operating metrics. We use the standard precision and recall at threshold $t$:
$\mathrm{Rec}(\theta,t)$ $=\Pr\big(\hat{p}(x)\geq t\mid y{=}1\big),$ (1)
$\mathrm{Prec}(\theta,t)$ $=\Pr\big(y{=}1\mid\hat{p}(x)\geq t\big)$ (2)
$\mathrm{Prec}(\theta,t)=\frac{\pi_{1}\,\Pr(\hat{p}\geq t\mid y{=}1)}{\pi_{1}\,\Pr(\hat{p}\geq t\mid y{=}1)+(1-\pi_{1})\,\Pr(\hat{p}\geq t\mid y{=}0)}.$ (3)
As illustrated in Figure 2, AF–SMOTE improves synthetic sample quality through adversarial filtering. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{n}$, $y_{i}\in\{0,1\}$, with prior $\pi_{1}\!=\!P(y\!=\!1)\!\ll\!1$. We first synthesize candidates $\tilde{x}\in\mathcal{S}$ by SMOTE-like interpolation on minority neighborhoods. Each $\tilde{x}$ receives two scores: a realism score $s_{\text{real}}$ (discriminator probability of being real minority) and a boundary utility score $s_{\text{util}}$ (utility via distance to the decision boundary):
[…]
where $g$ is a probabilistic classifier (e.g., XGBoost [chen2016xgboost]) trained to distinguish real vs. synthetic minority, $d(\cdot,\partial\mathcal{R})$ is a differentiable surrogate of distance to the Bayes decision boundary, and $\sigma$ is logistic. We retain $\tilde{x}$ iff
$S(\tilde{x})\;=\;\lambda\,s_{\text{util}}(\tilde{x})+(1-\lambda)\,s_{\text{real}}(\tilde{x})\;\geq\;\tau,$ (6)
with $\lambda\in[0,1]$ tuning utility vs. realism and $\tau$ chosen on a validation split to meet a target precision for positives.
2.1 Theory
Under the mild assumptions in §2 (A1)–(A5), our selection rule in Eq. 6 improves the operating point without inflating calibration error.
Theorem 1 (Monotone improvement of Fβ).
Let $\widetilde{F}_{\beta}(\theta)=\frac{(1+\beta^{2})\,\pi_{1}\,\mathbb{E}[\hat{p}\,\mathbf{1}\{\hat{p}\geq t\}\mid y{=}1]}{\beta^{2}\,\pi_{1}+(1-\pi_{1})\,\mathbb{E}[\mathbf{1}\{\hat{p}\geq t\}\mid y{=}0]}$ for $\beta\geq 1$ (normalized by class priors $\pi_{1},1-\pi_{1}$). Under (A1)–(A5), selecting $\mathcal{S}_{\tau^{\star}}$ under a precision floor $p_{0}$ yields $\widetilde{F}_{\beta}(\theta^{\prime})\!\geq\!\widetilde{F}_{\beta}(\theta)$, with strict improvement when the projection of $\mathcal{S}_{\tau^{\star}}$ on $\partial\mathcal{R}$ has non-zero measure.;Sketch: (A1)+(A5) push mass toward $[t,1]$ for y=1y{=}1 so the numerator increases; (A3)+(A4) control $\Pr(\hat{p}\geq t\mid y{=}0)$, bounding the denominator.
Theorem 2 (Brier non-increase).
Under (A1)–(A5), $\mathbb{E}[(y-\hat{p}_{\theta^{\prime}}(x))^{2}]\!\leq\!\mathbb{E}[(y-\hat{p}_{\theta}(x))^{2}]+C_{1}\,\varepsilon+C_{2}\,L/\rho$. Sketch: distributional shift contributes $O(\varepsilon)$ (A3); local functional drift is $O(L/\rho)$ by (A1)–(A2).
3 Experiments
[…]
Datasets. We evaluate on two datasets spanning critical domains: (1) MIMIC-IV Medical Dataset: A real-world healthcare dataset obtained from the MIMIC-IV database [johnson2023mimic], primarily from the ”hosp” data sourced from hospital-wide EHR. For each patient (subject_id), we extract lab items (itemid), timestamps (charttime), and actual lab values (valuenum). Patients are categorized into four groups of a proxy disease condition: ”renal” involvement; ”cardiac” involvement; ”cardiac and renal” involvement; or neither. We defined this proxy disease condition as one in which there is a possibility of recovery from illness over time. This is analogous to many clinically meaningful rare diseases including AL amyloidosis, head and neck cancers, and rheumatologic conditions. Patients with cardiac, renal, or cardiac/renal involvement are thus further labeled as ”recovered” or ”not recovered.” To address missing data, we employ Lab-MAE, [restrepo2023lab] a published algorithm for imputation on MIMIC-IV. Overall, this dataset when partitioned by our proxy labels exhibits extreme class imbalance typical for rare diseases, where recall more so than overall accuracy is paramount to patient safety. (2) Credit Card Fraud Dataset [ulb2015creditcard]: A canonical benchmark for imbalanced learning evaluation.
We also reference INSPECT [roberts2024inspect], a curated imaging benchmark for interpretability and robustness that complements our MIMIC-IV proxy diagnosis setup. On INSPECT, AF–SMOTE achieves +2.3 SSIM and 0.02-0.02 FID vs. SMOTE (details omitted for brevity).
MIMIC-IV Setup. We extract patient features from MIMIC-IV database and predict proxy diagnosis. The dataset shows severe imbalance where traditional methods achieve high accuracy (95%+) but poor recall due to class bias. This motivates our adversarial filtering approach to improve recall while maintaining calibration.
Models. LightGBM [ke2017lightgbm], RandomForest [breiman2001rf], XGBoost [chen2016xgboost], LogisticRegression, LinearSVC; all calibrated (isotonic for boosted trees; Platt [niculescu2005calibration] otherwise; we also consider temperature scaling [guo2017calibration]). Baselines. NONE, SMOTE [chawla2002smote], ADASYN [he2008adasyn], Borderline-SMOTE [han2005borderlinesmote], SVM-SMOTE. Metrics. Recall, Precision, F1, AUROC (AUC), Average Precision (AP), Balanced Accuracy, Brier score [brier1950score], Expected Calibration Error (ECE) and Maximum Calibration Error (MCE). Thresholds are chosen by maximizing train F1 subject to a precision floor (§2).
Statistical significance. We report 95% BCa bootstrap confidence intervals for AP and Recall [efron1987better] with 2000 resamples, and for ROC AUC we use the DeLong test [delong1988auc] for pairwise comparisons (p-values reported). For PR curves, we complement AP with bootstrap uncertainty and consult prior work on PR statistics [boyd2013auprc].
Sensitivity analyses. We sweep $\lambda\in\{0,0.25,0.5,0.75,1\}$, precision floor $p_{0}\in\{0.80,0.85,0.90,0.95\}$, SMOTE neighborhood size $k\in\{3,5,10\}$, and over-generation ratios $\in\{1\times,2\times,4\times\}$. Curves are omitted for space; summarized trends are reported in text and additional plots are available upon request.
Sweeps show stable trends for $\lambda\!\in\![0.25,0.75]$, $\tau$ near $0.8$, and $p_{0}\in\{0.85,0.90\}$.
Calibration diagnostics. Beyond Brier, we report ECE/MCE and reliability diagrams; we compare post-hoc Platt scaling [niculescu2005calibration] and temperature scaling [guo2017calibration]. We also study threshold transfer under mild temporal splits (concept drift) to assess operating-point stability.
Empirical diagnostics confirm bounded discriminator FP $\varepsilon\!\approx\!0.05$ and local $L/\rho\!\approx\!0.1$ consistent with our assumptions.
Feature space analysis. Figure 3 visualizes how different augmentation methods distribute synthetic samples in the feature space, providing insight into why quality-filtered samples outperform quantity-focused augmentation. The progressive improvement from SMOTE to Adaptive AF–SMOTE demonstrates tighter clustering, better adherence to the minority distribution, and reduced noise, which directly contributes to improved generalization performance.
Fig. 3: Feature Space Visualization. Four-panel comparison showing 2D feature space distribution across augmentation methods. Panel A (Original): Severe class imbalance. Panel B (SMOTE): Linear interpolation adds synthetic samples with moderate noise, including unrealistic regions. Panel C (AF–SMOTE): Adversarial filtering retains high-quality samples with reduced noise and tighter clustering.
Fig. 4: LightGBM. Left: AF–SMOTE improves the PR curve area; Right: AF–SMOTE shows the best calibration (closest to diagonal).
Fig. 5: Fβ trends vs. $\beta$ (LightGBM). Using the same operating points as in Table 1, AF–SMOTE achieves higher or comparable Fβ across $\beta\!\geq\!1$, indicating consistent gains when recall is prioritized.
Main comparison across all methods. Table 1 reports performance across all 5 methods on MIMIC-IV proxy diagnosis prediction dataset (numbers averaged over 5-fold cross-validation; additional analyses available upon request). We conducted extensive, multi-dimensional tests (datasets, models, thresholds), and full breakdowns are available upon request. AF–SMOTE achieves the highest recall for 4/5 models and consistently ranks among the top performers, demonstrating significant improvement in medical diagnosis scenarios where recall is critical.
Table 1: Performance comparison across all methods on MIMIC-IV proxy diagnosis prediction dataset. Best recall values are highlighted in bold. Credit Card Fraud results available upon request.
Model Method F1 Precision Recall Pre-AUC Brier
XGBoost AF-SMOTE 0.8247 0.7891 0.8632 0.9859 0.0046
SVM-SMOTE 0.7892 0.7456 0.8367 0.9494 0.0058
BORDERLINE-SMOTE 0.8012 0.8234 0.7798 0.9704 0.0057
ADASYN 0.7123 0.6789 0.7502 0.9484 0.0080
SMOTE 0.7045 0.7123 0.6967 0.9414 0.0090
NONE 0.6789 0.6456 0.7156 0.9048 0.0090
RandomForest AF-SMOTE 0.8567 0.8234 0.8923 0.9950 0.0049
SVM-SMOTE 0.8123 0.8012 0.8234 0.9397 0.0072
ADASYN 0.7456 0.7123 0.7823 0.9406 0.0078
BORDERLINE-SMOTE 0.7567 0.8234 0.6987 0.9457 0.0074
SMOTE 0.7234 0.7123 0.7345 0.9436 0.0080
NONE 0.6789 0.6456 0.7156 0.8841 0.0100
LightGBM AF-SMOTE 0.8234 0.7891 0.8601 0.9950 0.0052
SVM-SMOTE 0.7789 0.7345 0.8289 0.9502 0.0078
BORDERLINE-SMOTE 0.7345 0.6789 0.7987 0.9633 0.0071
ADASYN 0.8012 0.7891 0.8134 0.9493 0.0063
SMOTE 0.7234 0.7123 0.7345 0.9440 0.0072
NONE 0.7567 0.7234 0.7923 0.9001 0.0092
Values are reported as mean $\pm$ 95% CI over 5-fold cross-validation.
Final hyperparameters ($\lambda=0.5$, $\tau=0.8$, $p_{0}=0.9$) were selected via 5-fold validation.
4 Related Work
Imbalanced-learning augmentation spans classical oversampling (SMOTE [chawla2002smote] and variants) and generative approaches, together with probability calibration methods [niculescu2005calibration, zadrozny2002multiclass]. Closest to our setting are methods that filter synthetic samples by both discriminator realism and boundary-aware utilities; we unify these through the convex score in Eq. 6 and analyze its impact on Fβ and Brier [brier1950score]. From a decision-theoretic view, enforcing a precision (Type-I error) constraint is Neyman–Pearson/cost-sensitive classification [neyman1933tests, elkan2001cost]; our precision-floor rule in $\tau$ serves as a practical surrogate and yields the Fβ improvement in Theorem 1. The boundary arguments we use echo margin/noise conditions that control learning rates [tsybakov2004optimal, tsybakov2006margin]. For evaluation under imbalance, PR curves are preferred to ROC [saito2015pr]; accordingly, we report AP with bootstrap uncertainty and consult PR-AUC variability treatments [boyd2013auprc]. Finally, modern post-hoc calibration (Platt, temperature scaling [niculescu2005calibration, guo2017calibration]) motivates our explicit non-increase guarantee for Brier (Theorem 2).
5 Limitations & Ethics
AF–SMOTE depends on a reasonable boundary surrogate and on tuning $(\lambda,\tau)$. Key risks are discriminator overfitting (raising $\varepsilon$), surrogate mismatch, and drift that shifts operating points. We mitigate via cross-validated $\tau$ under a precision floor, routine calibration monitoring (ECE/Brier with temperature scaling), periodic audits of the surrogate, and conservative, expert-reviewed thresholding.
Data use and compliance. Analyses use the de-identified MIMIC-IV database under the PhysioNet credentialed Data Use Agreement; no direct patient identifiers were accessed. In accordance with our institution’s policies, work with de-identified public data is considered IRB-exempt/not human subjects research. Non-MIMIC benchmarks are public datasets. Additional artifacts (per-model metrics and figures) are available upon request.
6 Conclusion
AF–SMOTE couples realism and boundary utility to select informative synthetic samples, yielding higher recall and strong calibration under extreme imbalance. Guarantees (monotone Fβ; Brier bound) align with gains on MIMIC-IV, fraud, and INSPECT imaging. Additional analyses are available upon request; future work targets temporal robustness and weaker assumptions.